Greetings!
Okay, who can help me understand word problems? I really need to know how to break down this problem and solve it. Could someone please help understand!
Problem: Alonzo travels 230 miles in 3 hours, how many miles miles traveled in 7 hours?
Thanks to all.

Answer by venugopalramana(3286)

Alonzo travels 230 miles in 3 hours, how many miles miles traveled in 7 hours?
it is assumed that the rate of travel is uniform ..that is if he goes 100 miles in 1 hour ,then he will go 200 miles in 2 hours,300 miles in 3 hrs...etc....
so we get the eqn.
d=r*t..where d=distance travelled,r=rate or speed of travel and t=time..
let us first find his rate of travel.we have d=230..t=3...r=?
230=3r...or r=230/3
now to the second case we have r=230/3...t=7...d=?
d=7*230/3=1610/3=536.66666